Let y=3x^2.

Find the change in y, Δy when x=4x and Δx=0.2


Find the differential dy when x=4x and dx=0.2
Answer #1
  1. Anu 3 X x-4 amd bx-os in the above equation oy = 9.930 4Oplease like
Answer #2

Let y=2x
Find the change in yΔy when x=4 and Δx=0.2
Answer #3

Our first order of business here is to find the derivative of the equation (i.e. differential equation). This requires u-substitution...

 

u = 5x + 6

du = 5 dx

 

Now to the equation...

 

y = tan u

dy = sec2u * du

     = sec2(5x+6)*5dx

 

So now we have our differential equation. Onward...

 

Solve for dy when x = 4 and dx = 0.4...

 

dy = sec2(5*4 + 6)*5*0.4

dy = sec2(26) * 2

dy = 2.38946 * 2

dy = 4.77892

 

Yay! The second case is solved the same way. :)


answered by: zahid